larry is 2 years younger then Ben. if the sum of their age is 49, what are their ages?
Answer by jennywecas(28)

L=larry
B=Ben
L=B-2
L+B=49
B-2+B=49
2B=51
b=25.5
Bill is 25.5 years old and Larry 23.5 years old